OpenStudy (anonymous):

ill medal and fan: Reiko started a business selling home medical supplies. She spent $5200 to obtain her merchandise, and it costs her $550 per week for general expenses. She earns $900 per week in sales. What is the minimum number of weeks it will take for Reiko to make a profit? A. 15 weeks B. 14 weeks C. 13 weeks D. 12 weeks

OpenStudy (anonymous):

A. 15weeks because- (make)900*15=13,500 (spent)550*15=8250+5200=13,450 you make a 50$profit in15weeks